# reference.bib

### v1

@@ 2 The split conformal procedure
$1-\alpha\leq\mathbbm{P}(Y_{n+1}\in{C_{n}}(X_{n+1}))\leq 1-\alpha+\frac{2}{|I_{2}|+2},$
@@ 3 Boosted conformal procedure
On the training set, call GradientBoosting (10) to obtain boosted functions $\{E_{j}^{(t)}\}_{t=0}^{\tau}$
@@ 5.2.1 Theoretical guarantees
Theorem 5.1 (Asymptotic expressiveness).
@@ 5.2.2 Empirical results on real data
The experiment results summarized in Table 1 indicate that applying boosting significantly enhances the performance of the baseline Local procedure. In contrast, boosting on CQR does not yield significant improvements—a sign that CQR already targets conditional coverage. (Before boosting, the prediction intervals generated by the baseline Local procedure exhibit conditional coverage deviations up to three times greater than those of the baseline CQR procedure.) It is noteworthy, however, that after boosting, the conditional coverage of the Local procedure improves to a level comparable to that of the boosted CQR procedure. While generally slightly less effective, nevertheless surpasses the performance of the boosted CQR procedure in two cases. Results on the remaining datasets are deferred to Table A2.
@@
meps-21 17.627 23917.62723917.627\,23917.627 239 7.832 -$\text{-55.57}\%-$ 8.164 7358.1647358.164\,7358.164 735 8.067 -1.2%
6 Boosting for length/power
We begin by specifying the oracle prediction interval with minimum length. For a random variable $Z$, the High Density Region (HDR) at a specified significance level $\alpha$, denoted as $\text{HDR}_{\alpha}(Z)$, is defined as the shortest deterministic interval that covers $Z$ with probability at least $1-\alpha$. The boundaries of $\text{HDR}_{\alpha}(Z)$, the lower limit $Q_{l(\alpha)}$ and the upper limit $Q_{u(\alpha)}$, obey the condition $\mathbbm{P}(Z\in[Q_{l(\alpha)},Q_{u(\alpha)}])\geq 1-\alpha$. For a pair of $(X,Y)$ drawn from $P$, for every value of $x\in\mathbb{R}^{p}$, the most powerful oracle prediction [^3] We take the liberty of using the term power due to the analogy with statistical testing in which a more powerful test leads to shorter confidence intervals. interval at that point is expressed as
@@ 6 Boosting for length/power
Before introducing the strategy to boost power, we present a word of caution against optimizing exclusively for this objective. Importantly, to maintain valid marginal coverage, the most powerful prediction interval is prone to overcover when the spread of $Y|X$ (the conditional distribution of $Y$ given $X$) is small, and undercover when the spread of $Y|X$ is large. This may be undesirable.
Similar to Theorem 5.1, we can show that the generalized score families exhibit the necessary expressiveness to contain the oracle conformity score, achieving optimal length while ensuring valid marginal coverage. The formal theorem is deferred to Section A.3.
6.1 A measure for length/power
@@
$\ell_{P}({E};\mathcal{D})=n^{-1}{\textstyle\sum}_{i=1}^{n}|{C_{n}(}X_{i})|.$ (18)
To derive a differentiable approximation of ${\ell}_{P}$, we approximate the empirical quantile $Q_{1-\alpha}$ in the conformalized intervals (6) and (9) with the smooth quantile estimator ${Q^{s}_{1-\alpha}}$ constructed in (15). Here, we demonstrate calculating the derivative of the smooth approximation of ${\ell}_{P}$ with respect to each component of the generalized Local score, expanding it as follows based on the previously outlined approximation steps:
${\ell}_{P}\left({E};\mathcal{D}\right)=2\bar{\sigma}{Q}_{1-\alpha}^{s}(\{E_{i}\}_{i=1}^{n}),\qquad E_{i}=|Y_{i}-\mu_{i}|/\sigma_{i},$
with $\mu_{i}=\mu(X_{i})$, $\sigma_{i}=\sigma(X_{i})$, $\bar{\sigma}=n^{-1}\sum_{i=1}^{n}\sigma_{i}$. As a result, for each feature $X_{i}$ within $\mathcal{D}$, we can evaluate $\partial{\ell}_{P}\left({E};\mathcal{D}\right)/\partial\mu_{i}$ and $\partial{\ell}_{P}\left({E};\mathcal{D}\right)/\partial\sigma_{i}$ via the chain rule. For instance,
${\partial{\ell}_{P}\left({E};\mathcal{D}\right)}/{\partial\mu_{i}}=-2\bar{\sigma}[{\partial{Q}_{1-\alpha}^{s}(\{E_{j}\}_{j=1}^{n})}/{\partial E_{i}}]\cdot[{\text{sign}(Y_{i}-\mu_{i})}/{\sigma_{i}}].$
@@ 6.2 Empirical results on real data
We apply our boosted conformal procedure to the same datasets described in Section 5.2.2. Detailed information on the models and hyperparameters used during the training and boosting stages can be found in Section A.7. Partial experiment results are summarized in Table 2. Notably, the boosting performance highlighted in bold exhibits significant improvement compared to previously documented results [romano2019conformalized, Sesia_2020]. We see a pronounced enhancement with the blog dataset; before boosting, the Local prediction intervals are on average $47\%$ longer than those generated by CQR. After boosting, these intervals outperform the boosted CQR intervals by $20\%$. Using CQR as the baseline also yields substantial improvements, a decrease in averaged length exceeding $10\%$ in six out of the eleven datasets. The meps-21 dataset, in particular, shows an improvement of up to $17\%$ relative to the baseline. Results on the remaining datasets can be found in Table A3. Figure 4 compares the conformalized prediction intervals derived from baseline Local and CQR scores with those obtained from the boosted scores. To effectively visualize the impact of boosting, we conduct a regression tree analysis on the training set to predict the label $Y$, setting the maximum number of tree nodes to four. This regression tree is then applied to the test set, allowing for a detailed comparison of the prediction intervals across each of the four distinct leaves.
Figure 4: Comparison of test set power (averaged interval length) evaluated on the meps-19 and blog datasets: classical Local and CQR conformal procedure versus the boosted procedures (abbreviated as ‘Localb’ and ‘CQRb’) compared in each of the 4 leaves of a regression tree trained on the training set to predict the label $Y$. A positive log ratio value between the regular and boosted interval lengths indicates improvement from boosting. The target miscoverage rate is set at $\alpha=10\%$.
Table 2: Test set power $\ell_{P}$ evaluated on various conformalized prediction intervals. The best result achieved for each dataset is highlighted in bold.
Power Loss (Average Length), target miscoverage $\alpha=10\%$
@@
blog 2.056 2562.0562562.056\,2562.056 256 1.068 -$\text{-48.08}\%-$ 1.395 2511.3952511.395\,2511.395 251 1.382 -$\text{-0.93}\%-$
facebook-1 1.896 4651.8964651.896\,4651.896 465 1.443 -$\text{-18.76}\%-$ 1.184 5591.1845591.184\,5591.184 559 1.056 -$\text{-10.85}\%-$
facebook-2 1.854 0891.8540891.854\,0891.854 089 1.319 -$\text{-28.87}\%-$ 1.220 9061.2209061.220\,9061.220 906 1.055 -$\text{-13.57}\%-$
meps-19 2.070 4032.0704032.070\,4032.070 403 1.695 -$\text{-18.15}\%-$ 2.553 1172.5531172.553\,1172.553 117 2.205 -$\text{-13.64}\%-$
meps-20 2.081 1532.0811532.081\,1532.081 153 1.836 -$\text{-12.26}\%-$ 2.668 8772.6688772.668\,8772.668 877 2.346 -$\text{-12.09}\%-$
meps-21 2.062 9272.0629272.062\,9272.062 927 1.805 -$\text{-12.49}\%-$ 2.620 2992.6202992.620\,2992.620 299 2.173 -$\text{-17.06}\%-$
@@ 7 Discussion
We introduced a post-training conformity score boosting scheme aiming to optimize for conditional coverage and power of the conformalized prediction interval. An intriguing avenue for future exploration involves simultaneously optimizing both power and conditional coverage, potentially incorporating user-specified weights for these objectives. Additionally, we can readily adapt our procedure to meet various application-specific objectives. For instance, we can optimize for conditional coverage on predefined feature groups, a common task in enhancing fairness in distributing social resources across different demographic groups [gibbs2023conformal]. Similarly, we can modify our procedure to reduce the length of prediction intervals for predefined label groups, which can be seen as reallocating resources to decrease uncertainty for certain groups at the expense of higher uncertainty for other groups [stutz2022learning]. Candidate loss functions tailored to these objectives are detailed in Section A.1. Lastly, the primary emphasis of this paper centers on the design of the conformity score boosting scheme and formalizing the optimization of conditional coverage in mathematical terms, leaving room for computational optimization to enhance performance and runtime efficiency.
@@ A.3 Boosting for length: theoretical guarantees
Similar to Theorem 5.1, we show in Theorem A.2 below that the generalized Local and CQR score families exhibit the necessary expressiveness to contain the oracle score, achieving optimal length while ensuring valid marginal coverage.
Theorem A.2 (Asymptotic expressiveness).
Under the assumptions of Theorem 5.1, for any target coverage rate $1-\alpha$, as $n\rightarrow\infty$, the following statements hold true:
@@
    There exists $(\mu^{*},\sigma^{*})$ such that the corresponding generalized Local score function (5) recovers the most powerful oracle prediction interval (17).
@@
    There exists $(\mu_{1}^{*},\mu_{2}^{*},\sigma^{*})$ such that the corresponding generalized CQR score function (8) recovers the most powerful oracle prediction interval (17).
Proof of Theorem A.2.
@@ A.8 Additional results and error bars
In Tables A2 and A3, we report additional experiment results for the datasets not included in Tables 1 and 2.
Table A2: Test set maximum deviation loss $\ell_{M}$ evaluated on various conformalized intervals.
[…]
Dataset Method Improvement Method Improvement
Local Boosted CQR Boosted
blog 51.352 65651.35265651.352\,65651.352 656 48.591 08248.59108248.591\,08248.591 082 -$\text{-5.38}\%-$ 27.759 66527.75966527.759\,66527.759 665 26.836 19226.83619226.836\,19226.836 192 -3.33%
facebook-1 26.020 33926.02033926.020\,33926.020 339 25.917 05225.91705225.917\,05225.917 052 -$\text{-0.40}\%-$ 13.407 33613.40733613.407\,33613.407 336 13.255 36913.25536913.255\,36913.255 369 -1.13%
facebook-2 42.807 42742.80742742.807\,42742.807 427 42.436 50142.43650142.436\,50142.436 501 -$\text{-0.87}\%-$ 18.256 85118.25685118.256\,85118.256 851 18.002 42818.00242818.002\,42818.002 428 -1.39%
star 9.658 3689.6583689.658\,3689.658 368 9.348 4289.3484289.348\,4289.348 428 -$\text{-3.21}\%-$ 9.727 9379.7279379.727\,9379.727 937 9.630 3429.6303429.630\,3429.630 342 -1.0%
Table A3: Test set power $\ell_{P}$ evaluated on various conformalized prediction intervals.
Power Loss (Average Length), target miscoverage $\alpha=10\%$
Dataset Method Improvement Method Improvement
Local Boosted CQR Boosted
bike 1.774 8091.7748091.774\,8091.774 809 1.402 5381.4025381.402\,5381.402 538 -$\text{-20.98}\%-$ 0.554 5730.5545730.554\,5730.554 573 0.497 9020.4979020.497\,9020.497 902 -$\text{-10.22}\%-$
bio 1.561 8801.5618801.561\,8801.561 880 1.520 0151.5200151.520\,0151.520 015 -$\text{-2.68}\%-$ 1.517 3601.5173601.517\,3601.517 360 1.471 0531.4710531.471\,0531.471 053 -$\text{-3.05}\%-$
community 1.727 8381.7278381.727\,8381.727 838 1.718 4231.7184231.718\,4231.718 423 -$\text{-0.54}\%-$ 1.710 0391.7100391.710\,0391.710 039 1.694 1551.6941551.694\,1551.694 155 -$\text{-0.93}\%-$
concrete 1.010 3621.0103621.010\,3621.010 362 0.680 7170.6807170.680\,7170.680 717 -$\text{-32.63}\%-$ 0.391 9310.3919310.391\,9310.391 931 0.386 5060.3865060.386\,5060.386 506 -$\text{-1.38}\%-$
star 0.178 9140.1789140.178\,9140.178 914 0.176 8800.1768800.176\,8800.176 880 -$\text{-1.14}\%-$ 0.194 9600.1949600.194\,9600.194 960 0.194 5160.1945160.194\,5160.194 516 -$\text{-0.23}\%-$
@@
We have previously reported the evaluated losses $\ell_{M}$ and $\ell_{P}$ for each dataset, averaged over ten random seeds. Tables A4 and A5 below detail the distribution of these evaluations, providing the mean, $10\%$ quantile, and $90\%$ quantile for the test set deviations in conditional coverage ($\ell_{M}$) and power loss ($\ell_{P}$). These statistics are derived from 110 test set evaluations across 11 datasets and 10 random training-test splits. We opt to report empirical quantiles instead of standard deviations due to the asymmetric and non-Gaussian nature of the data.
@@
Table A5: Distribution of the test set power loss $\ell_{P}$ evaluated on various conformalized prediction intervals across 11 datasets and 10 random training-test splits.
Power Loss (Average Length), target miscoverage $\alpha=10\%$
@@
A.9 Additional figures on individual datasets

### v2

@@ 2 The split conformal procedure
$1-\alpha\leq\mathbbm{P}(Y_{n+1}\in{C_{n}}(X_{n+1}))\leq 1-\alpha+\frac{2}{|I_{2}|+2},$ (2)
@@ 3 Boosted conformal procedure
Since boosting is conducted on the training data, the boosted procedure satisfies the same marginal coverage guarantee as the split conformal procedure, as formalized below.
Proposition 3.1.
Let $\{(X_{i},Y_{i})\}_{i=1}^{m}$ be the held out calibration set, and $(X_{m+1},Y_{m+1})$ be a pair of new observation. If the $m+1$ samples are exchangeable, and ties between $\{E^{(\tau)}(X_{i},Y_{i})\}_{i=1}^{m}$ occur with probability zero, the confromalized prediction interval (1) computed from score function $E^{(\tau)}(\cdot,\cdot)$ satisfies the coverage guarantee (2).
@@
On the training set, call GradientBoosting (11) to obtain boosted functions $\{E^{(t)}\}_{t=0}^{\tau}$
@@ 5.2.1 Theoretical guarantees
Proposition 5.1 (Asymptotic expressiveness).
@@ 5.2.2 Empirical results on real data
The experiment results summarized in Table 1 indicate that applying boosting significantly enhances the performance of the baseline Local procedure. In contrast, boosting on CQR does not yield significant improvements—a sign that CQR already targets conditional coverage. (Before boosting, the prediction intervals generated by the baseline Local procedure exhibit conditional coverage deviations up to three times greater than those of the baseline CQR procedure.) It is noteworthy, however, that after boosting, the conditional coverage of the Local procedure improves to a level comparable to that of the boosted CQR procedure. While generally slightly less effective, nevertheless surpasses the performance of the boosted CQR procedure in two cases. Results on the remaining datasets are deferred to Tables A2 and A3.
@@
meps-21 17.627 23917.62723917.627\,23917.627 239 7.832 -$\text{-55.57}\%-$ 8.164 7358.1647358.164\,7358.164 735 8.067 -1.20%
6 Boosting for length
We begin by specifying the oracle prediction interval with minimum length. For a random variable $Z$, the High Density Region (HDR) at a specified significance level $\alpha$, denoted as $\text{HDR}_{\alpha}(Z)$, is defined as the shortest deterministic interval that covers $Z$ with probability at least $1-\alpha$. The boundaries of $\text{HDR}_{\alpha}(Z)$, the lower limit $Q_{l(\alpha)}$ and the upper limit $Q_{u(\alpha)}$, obey the condition $\mathbbm{P}(Z\in[Q_{l(\alpha)},Q_{u(\alpha)}])\geq 1-\alpha$. For a pair of $(X,Y)$ drawn from $P$, for every value of $x\in\mathbb{R}^{p}$, the oracle prediction
interval at that point is expressed as
@@ 6 Boosting for length
Before introducing the boosting strategy, we present a word of caution against optimizing exclusively for this objective. Importantly, to maintain valid marginal coverage, the shortest prediction interval is prone to overcover when the spread of $Y|X$ (the conditional distribution of $Y$ given $X$) is small, and undercover when the spread of $Y|X$ is large. This may be undesirable.
Similar to Proposition 5.1, we can show that the generalized score families
exhibit the necessary expressiveness to contain the oracle conformity score, achieving optimal length while ensuring valid marginal coverage. The formal proof is deferred to Section A.3.
6.1 A measure for length
@@
$\ell_{L}({E};\mathcal{D})=n^{-1}{\textstyle\sum}_{i=1}^{n}|{C_{n}(}X_{i})|.$ (19)
To derive a differentiable approximation of $\ell_{L}$, we approximate the empirical quantile $Q_{1-\alpha}$ in the conformalized intervals (7) and (10) with the smooth quantile estimator ${Q^{s}_{1-\alpha}}$ constructed in (16). Here, we demonstrate calculating the derivative of the smooth approximation of $\ell_{L}$ with respect to each component of the generalized Local score, expanding it as follows based on the previously outlined approximation steps:
$\ell_{L}\left({E};\mathcal{D}\right)=2\bar{\sigma}{Q}_{1-\alpha}^{s}(\{E_{i}\}_{i=1}^{n}),\qquad E_{i}=|Y_{i}-\mu_{i}|/\sigma_{i},$
with $\mu_{i}=\mu(X_{i})$, $\sigma_{i}=\sigma(X_{i})$, $\bar{\sigma}=n^{-1}\sum_{i=1}^{n}\sigma_{i}$. As a result, for each feature $X_{i}$ within $\mathcal{D}$, we can evaluate $\partial\ell_{L}\left({E};\mathcal{D}\right)/\partial\mu_{i}$ and $\partial\ell_{L}\left({E};\mathcal{D}\right)/\partial\sigma_{i}$ via the chain rule. For instance,
$\frac{\partial\ell_{L}\left({E};\mathcal{D}\right)}{\partial\mu_{i}}=-2\bar{\sigma}\frac{\partial{Q}_{1-\alpha}^{s}(\{E_{j}\}_{j=1}^{n})}{\partial E_{i}}\frac{\text{sign}(Y_{i}-\mu_{i})}{\sigma_{i}}.$
@@ 6.2 Empirical results on real data
We apply our boosted conformal procedure to the same datasets described in Section 5.2.2. Detailed information on the models and hyperparameters used during the training and boosting stages can be found in Section A.7. Partial experiment results are summarized in Table 2. Notably, the boosting performance highlighted in bold exhibits significant improvement compared to previously documented results [romano2019conformalized, Sesia_2020]. We see a pronounced enhancement with the blog dataset; before boosting, the Local prediction intervals are on average $42\%$ longer than those generated by CQR. After boosting, these intervals outperform the boosted CQR intervals by $32\%$. Using CQR as the baseline also yields substantial improvements, a decrease in averaged length exceeding $10\%$ in six out of the eleven datasets. The meps-21 dataset, in particular, shows an improvement of up to $18\%$ relative to the baseline. Results on the remaining datasets can be found in Tables A4 and A5. Figure 4 compares the conformalized prediction intervals derived from baseline Local and CQR scores with those obtained from the boosted scores. To effectively visualize the impact of boosting, we conduct a regression tree analysis on the training set to predict the label $Y$, setting the maximum number of tree nodes to four. This regression tree is then applied to the test set, allowing for a detailed comparison of the prediction intervals across each of the four distinct leaves.
Figure 4: Comparison of test set average interval length evaluated on the meps-19 and blog datasets: classical Local and CQR conformal procedure versus the boosted procedures (abbreviated as ‘Localb’ and ‘CQRb’) compared in each of the 4 leaves of a regression tree trained on the training set to predict the label $Y$. A positive log ratio value between the regular and boosted interval lengths indicates improvement from boosting. The target miscoverage rate is set at $\alpha=10\%$.
Table 2: Test set average interval length $\ell_{L}$ evaluated on various conformalized prediction intervals. The best result achieved for each dataset is highlighted in bold.
Average Length, target miscoverage $\alpha=10\%$
@@
blog 2.056 2562.0562562.056\,2562.056 256 0.972 -$\text{-52.74}\%-$ 1.444 6621.4446621.444\,6621.444 662 1.434 -0.71%
facebook-1 1.896 4651.8964651.896\,4651.896 465 1.383 -$\text{-27.03}\%-$ 1.198 1841.1981841.198\,1841.198 184 1.072 -10.47%
facebook-2 1.854 0891.8540891.854\,0891.854 089 1.363 -$\text{-26.51}\%-$ 1.199 7861.1997861.199\,7861.199 786 1.075 -10.41%
meps-19 2.070 4032.0704032.070\,4032.070 403 1.685 -$\text{-18.60}\%-$ 2.553 5762.5535762.553\,5762.553 576 2.136 -$\text{-16.35}\%-$
meps-20 2.081 1532.0811532.081\,1532.081 153 1.836 -$\text{-11.80}\%-$ 2.667 1302.6671302.667\,1302.667 130 2.357 -$\text{-11.62}\%-$
meps-21 2.062 9272.0629272.062\,9272.062 927 1.795 -$\text{-12.99}\%-$ 2.584 9832.5849832.584\,9832.584 983 2.105 -$\text{-18.55}\%-$
@@ 7 Discussion
We introduced a post-training conformity score boosting scheme aiming to optimize for conditional coverage or length of the conformalized prediction interval. An intriguing avenue for future exploration involves simultaneously optimizing both length and conditional coverage, potentially trading off these objectives by incorporating user-specified weights [dabah2024temperature]. Additionally, we can readily adapt our procedure to meet various application-specific objectives. For instance, we can optimize for conditional coverage on predefined feature groups, a common task in enhancing fairness in distributing social resources across different demographic groups [gibbs2023conformal]. Similarly, we can modify our procedure to reduce the length of prediction intervals for predefined label groups, which can be seen as reallocating resources to decrease uncertainty for certain groups at the expense of higher uncertainty for other groups [stutz2022learning]. Candidate loss functions tailored to these objectives are detailed in Section A.1. Lastly, the primary emphasis of this paper centers on the design of the conformity score boosting scheme and formalizing the optimization of conditional coverage in mathematical terms, leaving room for computational optimization to enhance performance and runtime efficiency. In essence, the gradient boosting algorithm in our procedure can be replaced with any gradient-based machine learning model. Thus, another interesting future direction would be to explore whether alternative algorithms could enhance performance.
Acknowlegements
E.C. was supported by the Office of Naval Research under Grant No. N00014-24-1-2305, the National Science Foundation under Grant No. DMS2032014, and the Simons Foundation under Award 814641. R.F.B. was supported by the Office of Naval Research via grant N00014-20-1-2337, and by the National Science Foundation via grant DMS-2023109.
@@ A.3 Boosting for length: theoretical guarantees
Similar to Proposition 5.1, we show in Proposition A.2 below that the generalized Local and CQR score families exhibit the necessary expressiveness to contain the oracle score, achieving optimal length while ensuring valid marginal coverage.
Proposition A.2 (Asymptotic expressiveness).
Under the assumptions of Proposition 5.1, for any target coverage rate $1-\alpha$, as $n\rightarrow\infty$, the following statements hold true:
@@
    There exists $(\mu^{*},\sigma^{*})$ such that the corresponding generalized Local score function (6) recovers the shortest oracle prediction interval (18).
@@
    There exists $(\mu_{1}^{*},\mu_{2}^{*},\sigma^{*})$ such that the corresponding generalized CQR score function (9) recovers the shortest oracle prediction interval (18).
Proof of Proposition A.2.
@@ A.8 Additional results and error bars
In Tables A2 to A5, we present additional results on marginal coverage, maximum conditional coverage deviation ($\ell_{M}$), and average interval length ($\ell_{P}$) for each real dataset (including those not reported in Tables 1 and 2), both before and after boosting. Notably, in each case, boosting is applied to optimize either conditional coverage or average interval length. As a result, the non-targeted characteristic may or may not improve after boosting.
Table A2: Additional information on conformalized intervals obtained before and after boosting for conditional coverage with the Local conformity score as baseline. The target miscoverage rate is set to $\alpha=10\%$.
Dataset $\ell_{L}$ $\ell_{M}(\%)$ Marginal Cov.(%)
Local Boosted Local Boosted Local Boosted
bike 1.7751.775{1.775}1.775 2.2012.201{2.201}2.201 10.979 5.638 89.927 89.646
bio 1.602 1.614 5.303 4.862 90.024 90.093
blog 2.0802.080{2.080}2.080 3.403 51.353 48.591 89.995 90.040
community 1.824 10.759 25.755 13.466 90.376 89.549
concrete 1.058 1.062 10.740 8.763 90.583 91.359
facebook-1 1.896 4.790 26.020 25.917 90.201 90.056
facebook-2 1.881 2.273 42.807 42.437 89.966 89.989
meps-19 2.074 2.926 15.357 5.656 90.120 90.497
meps-20 2.102 2.778 16.939 6.998 89.963 90.031
meps-21 2.069 2.537 17.627 7.832 90.064 90.054
star 0.189 0.179 9.658 9.348 90.808 90.831
Table A3: Additional information on conformalized intervals obtained before and after boosting for conditional coverage with the CQR conformity score as baseline. The target miscoverage rate is set to $\alpha=10\%$.
Dataset $\ell_{L}$ $\ell_{M}$(%) Marginal Cov.(%)
CQR Boosted CQR Boosted CQR Boosted
bike 0.555 0.540 4.934 4.925 90.073 90.184
bio 1.518 1.515 5.069 4.700 89.841 89.853
blog 1.761 1.766 27.760 26.836 90.222 90.244
community 1.718 1.740 12.688 12.105 90.340 90.194
concrete 0.484 0.489 9.039 8.265 90.451 90.652
facebook-1 1.374 1.371 13.407 13.255 90.465 90.247
facebook-2 1.465 1.409 18.257 18.002 89.763 90.001
meps-19 2.784 2.784 5.507 5.507 90.257 90.257
meps-20 2.769 2.743 7.614 7.184 89.991 90.006
meps-21 2.834 2.815 8.16 8.067 90.169 90.067
star 0.199 0.209 9.728 9.630 91.085 91.339
Table A4: Additional information on conformalized intervals obtained before and after boosting for length with the Local conformity score as baseline. The target miscoverage rate is set to $\alpha=10\%$.
Dataset $\ell_{L}$ $\ell_{M}$(%) Marginal Cov.(%)
Local Boosted Local Boosted Local Boosted
bike 1.775 1.360 22.590 19.616 89.927 89.862
bio 1.562 1.514 5.995 5.791 89.937 89.962
blog 2.056 0.972 52.440 54.858 89.988 89.978
community 1.728 1.678 26.066 24.822 89.499 89.323
concrete 1.010 0.698 11.029 10.518 90.631 90.728
facebook-1 1.896 1.384 26.020 34.259 90.201 89.944
facebook-2 1.854 1.363 42.624 50.697 90.020 89.972
meps-19 2.070 1.685 18.626 14.623 90.054 90.070
meps-20 2.081 1.836 17.897 14.643 89.869 89.849
meps-21 2.063 1.795 18.795 13.324 89.914 89.920
star 0.179 0.179 9.976 9.407 90.901 90.577
Table A5: Additional information on conformalized intervals obtained before and after boosting for length with the CQR conformity score as baseline. The target miscoverage rate is set to $\alpha=10\%$.
Dataset $\ell_{L}$ $\ell_{M}$(%) Marginal Cov.(%)
CQR Boosted CQR Boosted CQR Boosted
bike 0.553 0.489 9.530 10.092 90.041 90.418
bio 1.516 1.468 5.408 5.265 89.670 89.880
blog 1.445 1.434 29.875 34.011 90.149 90.260
community 1.693 1.699 14.006 13.536 89.499 89.699
concrete 0.391 0.393 10.342 10.700 88.932 89.223
facebook-1 1.198 1.073 20.132 30.901 89.937 90.013
facebook-2 1.200 1.075 29.233 34.475 90.035 90.010
meps-19 2.554 2.136 10.357 11.285 90.228 90.399
meps-20 2.667 2.357 10.826 10.720 89.875 89.838
meps-21 2.585 2.105 10.968 10.565 89.946 89.863
star 0.195 0.194 9.982 10.142 91.455 91.432
@@
We have previously reported the evaluated losses $\ell_{M}$ and $\ell_{L}$ for each dataset, averaged over ten random seeds. Tables A6 and A7 below detail the distribution of these evaluations, providing the mean, $10\%$ quantile, and $90\%$ quantile for the test set deviations in conditional coverage ($\ell_{M}$) and average interval length ($\ell_{L}$). These statistics are derived from 110 test set evaluations across 11 datasets and 10 random training-test splits. We opt to report empirical quantiles instead of standard deviations due to the asymmetric and non-Gaussian nature of the data.
@@
Table A7: Distribution of the test set average interval length $\ell_{L}$ evaluated on various conformalized prediction intervals across 11 datasets and 10 random training-test splits.
Average Length, target miscoverage $\alpha=10\%$
@@
A.9 Experiments under different miscoverage rates
In Tables A8 and A9, we illustrate the performance of boosting for length with the Local conformity score as baseline with miscoverage rates set to $5\%$ and $20\%$, respectively.
Table A8: Additional information on conformalized intervals obtained before and after boosting for length with the Local conformity score as baseline. The target miscoverage rate is set to $\alpha=5\%$.
Dataset $\ell_{L}$ $\ell_{M}$(%) Marginal Cov.(%)
Local Boosted Improvement Local Boosted Local Boosted
bike 2.523 1.828 -27.54% 11.553 11.626 94.927 94.972
bio 1.881 1.801 -4.25% 5.013 4.694 94.857 94.802
blog 4.217 1.923 -54.39% 37.976 34.505 95.027 95.048
community 2.578 2.217 -13.99% 16.406 14.554 95.414 94.812
concrete 1.129 0.838 -25.80% 8.55 8.086 94.709 94.515
facebook-1 2.667 2.262 -15.18% 20.85 20.052 95.104 95.049
facebook-2 2.441 2.092 -14.30% 41.645 39.943 94.941 95.016
meps-19 3.618 2.825 -21.92% 10.515 9.178 95.005 95.008
meps-20 3.951 3.049 -22.84% 11.043 9.693 94.907 94.922
meps-21 4.030 3.033 -24.74% 8.918 8.910 95.061 95.026
star 0.207 0.207 -0.11% 6.57 6.906 95.358 95.289
Table A9: Additional information on conformalized intervals obtained before and after boosting for length with the Local conformity score as baseline. The target miscoverage rate is set to $\alpha=20\%$.
Dataset $\ell_{L}$ $\ell_{M}$(%) Marginal Cov.(%)
Local Boosted Improvement Local Boosted Local Boosted
bike $1.312$ $0.983$25.07% $24.649$ $28.561$ $79.403$ $80.152$
bio $1.248$ $1.213$2.80% $10.838$ $10.540$ $79.729$ $79.594$
blog $1.904$ $0.513$73.07% $45.813$ $64.360$ $79.756$ $79.920$
community $1.337$ $1.234$7.71% $25.047$ $28.831$ $79.674$ $79.674$
concrete $0.833$ $0.549$34.10% $13.831$ $20.177$ $80.728$ $80.825$
facebook-1 $1.624$ $0.747$54.02% $29.384$ $47.994$ $80.190$ $79.731$
facebook-2 $1.580$ $0.749$52.60% $36.488$ $62.070$ $79.890$ $79.930$
meps-19 $1.821$ $1.020$44.03% $20.561$ $24.763$ $80.326$ $80.013$
meps-20 $1.843$ $1.095$40.60% $19.137$ $24.194$ $79.684$ $79.823$
meps-21 $1.831$ $1.064$41.86% $18.227$ $23.172$ $80.674$ $80.057$
star $0.142$ $0.141$0.71% $15.420$ $14.517$ $80.647$ $80.370$
A.10 Training a gradient boosting algorithm with our custom loss functions
Our proposed boosted conformal procedure serves as a post-training step designed to refine the conformity score $E(\cdot,\cdot;f)$ obtained during model training. This procedure can leverage pre-trained models when available. In the absence of pre-trained models, we can alternatively train a gradient boosting algorithm directly using our custom loss functions. For example, in the context of the local score, we may initialize with $\mu^{(0)}=0$, $\sigma^{(0)}=1$ and then apply our boosted conformal procedure. As discussed in Section 7, this approach is flexible enough to replace gradient boosting with any gradient-based algorithm, such as neural networks, trained under our custom loss functions. This framework aligns with that of [stutz2022learning], which introduces a neural network trained with a custom loss function to minimize the average prediction set size in classification tasks.
In Figure A1, we compare the performance of the two approaches optimizing for average interval length, searching within the generalized Local score family $\mathcal{F}$ defined in (6). The primary distinction between the two procedures lies in the initialization: the first approach employs $\mu^{(0)}=0$, $\sigma^{(0)}=1$, while the second derives $\mu^{(0)}$ and $\sigma^{(0)}$ from a trained random forests model. We run the experiments on the meps-19 dataset and compare the performance across different splits of the training and calibration data. In this context, the percentage of training data refers to the proportion of training data within the combined training and calibration datasets. Our results indicate that cross-validation selects a greater number of boosting iterations when we directly train the gradient boosting algorithm, resulting in longer runtime. However, the average interval length and maximum conditional coverage deviation after boosting are notably smaller for the boosted conformal procedure we introduced in this paper.
Figure A1: Comparison of boosted interval length, marginal coverage, maximum conditional coverage deviation ($\ell_{M}$), and runtime between direct training of a gradient-based algorithm (red) and boosting on a pre-trained conformity score (blue).
A.11 Selecting optimal boosting rounds via hold-out validation set
In our boosted conformal procedure, we use cross-validation on the training set to determine the optimal number of boosting rounds, a process that can be time-consuming. An alternative approach is to hold out a fraction of the training set for validation. While more computationally efficient, this method introduces a trade-off: a smaller validation set can lead to greater variability in prediction intervals and model performance, whereas a larger validation set may reduce the effective training set size, potentially limiting the model’s performance. To explore this trade-off, we conduct experiments on the bike dataset, optimizing for prediction interval length. We compare performance across two settings: 5-fold cross-validation and a hold-out validation set, with the training-to-validation set ratio ranging from 1:1 to 8:1. For each setting, we run 100 experiments, recording the average boosted length, the standard deviation of boosted lengths, and the average runtime. The results are shown in Figure A2.
Figure A2: Comparison of the average boosted length, standard deviation of boosted length, and average runtime of the boosting procedure when selecting the optimal number of boosting rounds using 5-fold cross-validation versus a hold-out validation set of varying sizes. Experiments are conducted on the bike dataset, with a target miscoverage rate of $\alpha=10\%$.
A.12 Additional figures on individual datasets